Grade 11Physical Chemistry

500ml of propane was burnt wid 3000ml of oxygen the yield of gaseous product was just 50% of the expected. find the total volume of all gases at the end of the reaction
REPLY FAST PLEASE

rajkumar jain
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
The complete balanced equation for the given reaction is:
C3H8(g) + 5O2(g)  3CO2(g) + 4H2O(g)

This means that
Amount of oxygen required for 1 mole of propane = 5 moles
So, volume of oxygen required by 500 mL propane =  5 × 500   =   2500   m L

Available oxygen = 300 mL
Thus , oxygen is the limiting reagent.

Now, 2500 mL oxygen will produce CO2 3 × 500   =   1500   m L
So, 300 mL oxygen will produce CO2 1500     m L 2500   m L × 300   m L  = 180 mL

Similarly, 2500 mL O2 will produce H2O =  4 × 500   =   2000   m L
So, 300 mL O2 will produce H2O =  2000   m L 2500   m L × 300   m L   =   240   m L

Total volume of products that should be formed = 180 + 240 mL = 420 mL
​But, total volume of products produced = 50% of 420 mL
                                                              =  50 100 × 420   =   210   m L

Now, since O2 is the limiting reactant, thus it is present in lesser amount than required for propane. So, some amount of propane will also be left in the reaction mixture after reaction.
So, 2500 mL oxygen requires propane = 500 mL
Thus, 300 mL oxygen will react with propane =  500   m L 2500   m L × 300   m L   =   60   m L

So volume of propane left unreacted = 500 - 60 mL = 440 mL

Oxygen is completely consumed.
So,  total volume of all gases that will be there at the end of the reaction = 210 + 440 mL
​                                                                                                                  = 650 mL